### v3

# FrugalPrompt: Reducing Contextual Overhead in Large Language Models via Token Attribution

Abstract
Human communication heavily relies on laconism and inferential pragmatics, allowing listeners to successfully reconstruct rich meaning from sparse, telegraphic speech. In contrast, large language models (LLMs) owe much of their stellar performance to expansive input contexts, yet such verbosity inflates monetary costs, carbon footprint, and inference-time latency. This overhead manifests from the redundant low-utility tokens present in typical prompts, as only a fraction of tokens typically carries the majority of the semantic weight. Inspired by the aforementioned cognitive psycholinguistic processes, we address this inefficiency by introducing FrugalPrompt, a novel prompt compression framework for LLMs, which retains only the most semantically significant tokens. Leveraging two state-of-the-art token attribution methods, GlobEnc and DecompX, we assign salience scores to every token in an input sequence, rank them to retain the top-$k\%$ tokens, and obtain a sparse frugalized prompt. We establish the theoretical stability of our approach and provide strong empirical results across a suite of four NLP tasks to study the trade-off between the portion of retained tokens and performance. Experimental findings across retention settings reveal asymmetric performance patterns that suggest potential task contamination effects. We posit that our work contributes to a more nuanced understanding of LLM behavior in performance-efficiency trade-offs and delineates the boundary between tasks tolerant of contextual sparsity and those requiring exhaustive context.
FrugalPrompt: Reducing Contextual Overhead in Large
Language Models via Token Attribution
Syed Rifat Raiyan11, Md Farhan Ishmam22, Abdullah Al Imran33, Mohammad Ali Moni44 11Islamic University of Technology, Bangladesh  22University of Utah, USA 33University of Liverpool, UK  44University of Queensland, Australia rifatraiyan@iut-dhaka.edu, farhan.ishmam@utah.edu
1 Introduction
As large language models (LLMs) are now becoming general-purpose intelligence systems for NLP, they are supporting workflows that go beyond their initial zero-shot text generation: from coding systems capable of using tools to multi-step agentic systems solving multimodal problems Yang et al. (2024); Schmidgall et al. (2025). Yet the same empirical scaling laws that fuel these capabilities into trillion-parameter models Fedus et al. (2022); Hudson et al. (2023) also intensify a basic cognitive and computational tension. Human language is richly redundant [^1] We define redundancy as any verbiage that extends beyond the most succinct form of communication., while modern LLM systems being trained and conditioned on human text are constrained by the context length Wu et al. (2025), latency Wan et al. (2023), and energy costs Stojkovic et al. (2024) due to the expressiveness of human language.
A core observation of studying linguistics from the lens of information theory is that natural language is not information-minimal code Shannon (1951). Human utterances inherently contain around $80\%$ statistical redundancy, e.g., text is repeated for clarification, and function words support grammatically more than semantic context. Many tokens, therefore, contribute little incremental information once the local context is known. Psycholinguistic evidence similarly suggests that comprehenders can often recover intended meaning from “telegraphic” input, a property exploited in simplified registers and second-language communication Gibson et al. (2013); Klein and Perdue (1997). These findings suggest a cognitively motivated route to efficiency in NLP, and we ask: To what extent do LLMs require full linguistic explicitness, and to what extent can they reconstruct meaning from pragmatically sufficient cues?
We aim to answer these questions with a theoretically grounded work motivated by information-theoretic approaches in linguistics. Using SOTA token attribution techniques, we compute saliency scores for the prompt and selectively retain the semantically rich tokens while filtering redundant ones. This yields a training-free compressor that (i) better preserves meaning-critical content, (ii) systematically removes low-value discourse material and function words, and (iii) avoids the cost of an auxiliary large model. Importantly, our approach requires only a lightweight pretrained encoder, providing substantially lower overhead than prior compressor designs. By introducing a prompt compression strategy that uses saliency scores generated from state-of-the-art pre-trained token attribution methods Modarressi et al. (2022, 2023). These saliency scores can capture the semantically rich token while filtering the redundant ones. This is achieved without any training paradigm, with minimal parameter overhead of a single 110M BERT model Devlin et al. (2019), substantially lower than that of existing methods.
Our criteria for filtering are grounded in two linguistic principles underlying textual redundancy. First, there is an intrinsic redundancy to any form of text produced by humans Wit and Gillette (1999). For instance, in the text, “What are your plans for the weekend, by the way?”, the phrase “by the way” contributes no additional meaning to an LLM. Second, the English language has low-importance tokens that add little to no semantic weight, generally in the form of articles and auxiliary verbs. For instance, most of the meaning is conveyed if the statement “I have received a parcel.”, is simplified to “I received parcel.”, with the latter mimicking simplified English, often used by non-native speakers Shubert et al. (1995). Our filtering methods eliminate both forms of redundancy. Our key contributions are as follows:
  • We introduce a novel, training-free prompt frugalization strategy for LLMs that controllably filters low-importance tokens based on saliency scores from pre-trained encoders.
  • Our method is theoretically grounded with empirical results showing strong retention–performance trade-offs. Across four NLP tasks, a $20\%$ prompt reduction preserves performance for most models with negligible parameter overhead.
  • We reveal behavioral insights into how scaling performance relates to inference cost and potential task contamination in benchmarks.
2 Related Work
2.1 Redundancy in Language
Language, by nature, is not information-minimal code Shannon (1951) and introduces several forms of redundancy. Through the lens of information theory Shannon (1948), redundancy of a language can be defined as the difference between the maximum possible entropy of the language and its actual entropy. Theories in inferential pragmatics Richards (1990) and relevance-driven communication argue that interlocutors reconstruct implicit content by focusing on semantically informative cues rather than processing every token with equal weight. Such laconism is not an anomaly but a fundamental property of human communicative efficiency Levshina and Moran (2021).
2.2 LLM Efficiency
The current literature on LLM efficiency Wan et al. (2023) stems from earlier works on efficient transformers Tay et al. (2022). Inference-time efficiency can be achieved via model compression methods, such as quantization Dettmers et al. (2023b), parameter pruning Ma et al. (2023), low-rank decomposition Li et al. (2023a), and knowledge distillation Timiryasov and Tastet (2023). Training efficiency can be achieved during pre-training Liu et al. (2023); Yao et al. (2024) and fine-tuning, either through parameter-efficient Hu et al. (2022) or memory-efficient Dettmers et al. (2023a) approaches. Finally, architectural changes also improve efficiency, e.g., efficient attention Child et al. (2019); Ainslie et al. (2023) and state-space models Gu and Dao (2023).
Distinct from these methods, our work adopts a data-centric approach, where the input prompt is optimized rather than the model’s underlying weights or architecture Chang et al. (2024). In essence, we perform prompt compression, which minimizes input redundancy while retaining task-specific information. Li et al. (2023b); Jiang et al. (2023). Contrasting previous approaches that often use a paraphrasing model Chuang et al. (2024), which can lead to instability, ours takes a theoretically grounded approach by filtering the least salient tail-end tokens.
2.3 Token Attribution
The self-attention weights in transformers have been primarily explored as a proxy for token importance Htut et al. (2019); Clark (2019), yet they fail to align with the true attribution of tokens Jain and Wallace (2019). While including vector norms can partially mitigate this issue Kobayashi et al. (2020), it overlooks the intricacies of the non-attention components in a transformer. Instead, we rely on subsequent token attribution methods (see Appendix A): GlobEnc Modarressi et al. (2022), which quantifies global token attribution by aggregating layer-wise analyses across the entire encoder using attention rollout and vector norms, and DecompX Modarressi et al. (2023), which propagates locally decomposed token representations globally to incorporate all model components, including non-linear feed-forward networks. These methods were selected for their superior ability to capture comprehensive semantic contributions beyond raw attention weights, enabling identification of salient tokens with minimal computational overhead.
Figure 1: Overview of the FrugalPrompt pipeline. The input prompt with tokens $\langle t_{1},t_{2},\dots,t_{n}\rangle$ is passed to our task-specific token attribution module to generate the saliency scores for each token $\langle s_{1},s_{2},\dots,s_{n}\rangle$. The tokens are ordered based on their saliency in a permutation $\pi$ and the top $p=k\%*n$ tokens are filtered while maintaining the order of the text to form the reduced text $\langle t_{\pi(1)},t_{\pi(2)},\dots,t_{\pi(p)}\rangle$. The reduced text is inferred by a frozen LLM to generate the answer tokens $\langle a_{1},a_{2},\dots,a_{n}\rangle$.
3 Methodology
Figure 1 portrays an overview of our proposed pipeline. We obtain the saliency scores using the token attribution methods: GlobEnc Modarressi et al. (2022) and DecompX Modarressi et al. (2023), details discussed in Appendix A and employ frugalization via token filtration to generate a condensed yet semantically representative version of an input sample. This technique identifies and retains only the most salient tokens, thereby reducing input dimensionality while preserving the core meaning for the given task.
3.1 Ranking via Saliency Scores
We define a task-specific scoring function $\varphi_{\tau}\in[\text{GlobEnc,\ DecompX}]$ for a task $\tau$, which assigns a saliency score to each token in a given text, i.e.,
$\varphi_{\tau}:\mathcal{T}\rightarrow\mathbb{R}^{n},$
where $\mathcal{T}$ denotes the space of the token sequences and $n$ is the length of the sequence. For the input token sequence, $T=\langle t_{1},t_{2},\dots,t_{n}\rangle$, the scoring function produces a saliency vector,
$\mathbf{s}=\varphi_{\tau}(T)=\langle s_{1},s_{2},\dots,s_{n}\rangle,\quad s_{i}\in\mathbb{R},$
where each score $s_{i}$ reflects the relative contribution of token $t_{i}$ to the model’s output for the specified task. Tokens are then ranked in monotone decreasing order of their attribution scores to obtain a permutation, $\pi$, formed of the token indices, such that,
$s_{\pi(1)}\geq s_{\pi(2)}\geq\cdots\geq s_{\pi(n)}.$
3.2 Top-$k$ Filtering
To construct the frugalized representation, we retain only the top $p=\left\lceil\frac{k}{100}\cdot n\right\rceil$ tokens, where $k$ is the retention percentage. The selected indices form the sequence,
${S}_{k}=\langle\pi(1),\pi(2),\dots,\pi(p)\rangle.$
We reindex $S_{k}$ in monotone increasing order to produce ${S}_{k\uparrow}$. The frugalized sequence ${F}_{k}$ is produced by restoring the selected tokens to their original order, i.e.,
${F}_{k}=\langle t_{i}\lvert i\in{S}_{k\uparrow}\rangle$
ensuring that the temporal coherence of the sequence is preserved, while uninformative tokens are discarded.
[…]
${F}_{40}=\langle\texttt{``movie''},\texttt{``good''},\texttt{``liked''},\texttt{``much''}\rangle$
where the original order is preserved to maintain syntactic coherence. The frugalized output “movie good liked much” is thus created, discarding the less salient tokens (e.g., ‘‘The’’, ‘‘was’’, etc.) while retaining semantic essence.
Model Token Attribu- tion $k\%$ Tasks
CLS SUM QA RSN
Acc F1 BLEU R-1 R-2 R-L BERT METEOR Acc pass@1
[…]
60 0.942 0.942 0.015 0.220 0.064 0.181 0.870 0.319 0.754 0.243
50 0.921 0.921 0.013 0.213 0.057 0.175 0.868 0.303 0.716 0.146
[…]
60 0.900 0.900 0.014 0.206 0.055 0.169 0.866 0.286 0.681 0.425
50 0.868 0.868 0.012 0.198 0.048 0.161 0.864 0.270 0.670 0.265
[…]
60 0.943 0.943 0.017 0.231 0.068 0.192 0.874 0.329 0.838 0.362
50 0.938 0.938 0.016 0.228 0.065 0.189 0.873 0.316 0.816 0.231
[…]
60 0.891 0.891 0.015 0.215 0.057 0.176 0.868 0.292 0.797 0.587
50 0.839 0.837 0.014 0.209 0.054 0.171 0.866 0.281 0.770 0.409
[…]
60 0.925 0.925 0.015 0.172 0.052 0.141 0.874 0.285 0.716 0.264
50 0.918 0.918 0.014 0.170 0.049 0.140 0.873 0.278 0.705 0.158
[…]
60 0.724 0.704 0.031 0.253 0.073 0.210 0.885 0.311 0.648 0.419
50 0.642 0.595 0.027 0.241 0.065 0.200 0.882 0.291 0.619 0.288
[…]
60 0.934 0.934 0.029 0.247 0.077 0.208 0.881 0.335 0.846 0.423
50 0.920 0.919 0.026 0.239 0.071 0.199 0.879 0.322 0.827 0.277
[…]
60 0.713 0.690 0.028 0.236 0.068 0.194 0.878 0.307 0.795 0.665
50 0.627 0.571 0.024 0.225 0.059 0.185 0.876 0.283 0.774 0.463
[…]
60 0.941 0.941 0.019 0.212 0.059 0.173 0.858 0.282 0.802 0.462
50 0.932 0.932 0.018 0.204 0.055 0.166 0.857 0.272 0.785 0.332
DecompX 80 0.842 0.839 0.020 0.208 0.056 0.170 0.858 0.273 $7 \cellcolormycyan0.8$50
60 0.727 0.707 0.018 0.195 0.049 0.159 0.854 0.253 0.724 0.679
50 0.641 0.593 0.017 0.187 0.045 0.152 0.853 0.236 0.686 0.533
Table 1: Impact of the two variants of FrugalPrompt retaining $k\%$ tokens on baseline LLM performance in text classification (CLS), text summarization (SUM), question answering (QA), and reasoning (RSN). The base and best prompt-reduced performance for each model are given in Gray and Light Blue cells, respectively. Bold and Underline indicate the highest overall performance and overall prompt-reduced performance, respectively.
3.3 Theoretical Bounds and Relationships
We derive the theoretical bounds of token filtering or deletion based on saliency scores. The bounds formalize the relationship between dropping $k\%$ tokens and the estimated performance.
Deletion Effect.
For a filtered or deleted set of indices $D\subseteq[n]=\{1,\dots,n\}$, we define the retained subsequence of tokens $T\setminus D:=T_{[n]\setminus D}$. For task $\tau$ and model parameters $\theta$, let $\ell_{\tau}(T,y;\theta)\in\mathbb{R}^{+}$ be the task loss or evaluation score. We define the deletion-effect set function:
$f(D)\;:=\;\underbrace{\ell_{\tau}(T\setminus D,y;\theta)}_{\text{Loss after Deletion}}-\underbrace{\ell_{\tau}(T,y;\theta)}_{\text{Original Loss}}.$ (1)
Trivially, $f(\emptyset)=0$. Eq. 1 is intuitively straightforward; $f(D)>0$ means deletion worsens the performance, whereas $f(D)<0$ means deletion improves performance, which is plausible. We emphasize that all results are inference-time, i.e., the model parameters $\theta$ are fixed. We use $\ell_{T}(T,y;\theta)$ solely as an evaluation functional to quantify how token deletion affects predictive performance.
Discrete Marginals and Interactions.
For a set of deleted tokens $A=[n]=\subseteq\{1,\dots,n\}$ and a new token $i\notin A$, we define marginal deletion:
$\Delta_{i}f(A)\;:=\;f(A\cup\{i\})-f(A).$ (2)
For tokens $i\neq j$ with $i,j\notin A$, we define the second-order discrete difference, i.e., the pairwise interaction as
$\Delta_{ij}f(A)\;:=\;\Delta_{i}\Delta_{j}f(A)=f(A\cup\{i,j\})$
$-f(A\cup\{i\})-f(A\cup\{j\})+f(A).$ (3)
Assumption 1 (Saliency upper-bounds singleton deletion harm).
There exists a task constant, $C_{\tau}\geq 0$, such that,
$f(\{i\})\;$ $\leq\;C_{\tau}\,s_{i},\quad\forall i\in[n].$ (4)
Assumption 1 is essentially true, as deleting token $i$ alone cannot increase the deletion effect more than a constant factor of its saliency.
Assumption 2 (Bounded pairwise interaction).
There exists a task constant, $\gamma_{\tau}\geq 0$, such that,
$\Delta_{ij}f(A)\;$ $\leq\;\gamma_{\tau},\quad\forall A\subseteq[n],\forall i,j\notin A.$ (5)
Assumption 2 simply imposes a pairwise bound that prevents the deletion effect of multiple tokens from causing the harm to explode.
Theorem 1 (Deletion bound via saliency and interaction).
Under Assumption 1, 2, for any deletion set $D$ with $|D|=q$,
$f(D)\;\leq\;C_{\tau}\sum_{i\in D}s_{i}\;+\;\frac{\gamma_{\tau}}{2}\,q(q-1).$ (6)
The proof of Theorem 1 is provided in Appendix §B. Theorem 1 establishes the stability of token deletion by upper-bounding the deletion effect.
Retention Relationship.
Following §3.2, we define the deleted set $D_{k}:=\{\pi(p+1),\dots,\pi(n)\}$ and the frugalized sequence $F_{k}:=T\setminus D_{k}$.
Corollary 1 (Performance vs. $k\%$ retention).
Under Assumption 1, with q=npq=n-p,
$\ell_{\tau}(F_{k},y;\theta)-\ell_{\tau}$ $(T,y;\theta)\;\leq\;C_{\tau}\sum_{j=p+1}^{n}s_{\pi(j)}\;$
$+\;\frac{\gamma_{\tau}}{2}\,(n-p)(n-p-1).$ (7)
Corollary 1 can be derived from Eq. 1, 5, and can also be generalized at the dataset level by computing the risk. Given $(T,y)\sim\mathcal{D}_{\tau}$ denotes the test distribution for task $\tau$, the risk can be simply defined as the expectation over the loss:
$R_{\tau}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\tau}}[\ell_{\tau}(x,y;\theta)]$ (8)
Figure 2: The performance difference between the reduced tokens and baseline performance using $100\%$ tokens of metrics across three tasks: text classification (CLS), question-answering (QA), and reasoning (RSN).
4 Experiment
With the performance bounds and relationship with the retention kk% are strongly conveyed in §3.3, we provide empirical evidence in this section.
4.1 Tasks
We evaluate the efficacy of the token reduction by experimenting across four NLP tasks: sentiment analysis (text classification), summarization (text generation), commonsense question answering, and mathematical reasoning. These tasks span both discriminative and generative paradigms, enabling evaluation under diverse linguistic and cognitive demands of the LLM.
4.2 Dataset
For sentiment analysis, we use IMDb movie review Maas et al. (2011), a binary sentiment analysis corpus comprising movie reviews labeled positive or negative. We use Argilla News to evaluate abstractive summarization on news articles Ahmed et al. (2017). For commonsense QA, we use CosmosQA Huang et al. (2019), which primarily focuses on commonsensical and causal reasoning within short narrative passages. The questions stipulated in CosmosQA are easy for humans to solve, but usually challenging for LLMs. We use GSM8k Cobbe et al. (2021) for mathematical reasoning, which contains grade-school MWPs to assess basic arithmetic reasoning and multi-step logical deduction.
Figure 3: The performance difference between the reduced tokens and baseline performance using $100\%$ tokens across the summarization metrics.
4.3 Models
We perform inference using a set of contemporary frontier models, varying in architecture and scale, to examine the generalizability of our method. The models are: LLaMA-3 8B, LLaMA-3 70B Meta (2024), GPT-3.5 Brown et al. (2020), Gemini 2.0 Flash Thinking Google DeepMind (2025), and o3-mini OpenAI (2025). The closed-source models are accessed via their respective APIs, and the open-source models are accessed via OpenRouter [^2] https://openrouter.ai/, using standardized prompts to minimize formatting-induced variability.
4.4 Experimental Setup
We set the temperature to 0 for all inference runs to sample from the model’s unaltered probability distribution, ensuring a standard and reproducible generation process. For the frugalized setup, we experiment by filtering the top $k\in\{50,60,80\}\%$ of the input tokens. We also include the baseline prompt, using $100\%$ of the tokens to set a proper frame of reference for the performance delta across all models. This setup enables us to investigate whether low-attribution filler tokens can be excluded with minimal degradation in task performance, thereby reducing API costs and carbon footprint.
4.5 Evaluation Metrics
We employ task-specific metrics to evaluate performance across both full and frugalized setups. For sentiment classification, we report accuracy and F1 score. Summarization performance is assessed using BLEU, ROUGE-1, ROUGE-2, ROUGE-L, BERT Score, and METEOR to capture both superficial and semantic quality of the responses. For commonsense question answering, we use accuracy as the dataset follows a multiple-choice format. Mathematical reasoning is evaluated using pass@$1$, which measures the proportion of problems correctly solved on the first attempt.
Attribution Type Token Attribution Method $k\%$ Tasks
CLS SUM QA RSN
o3-mini GPT-3.5 Gemini-2.0FT o3-mini
Acc F1 BLEU R-1 R-2 R-L BERT METEOR Acc pass@1
[…]
60 0.893 0.893 0.022 0.240 0.064 0.197 0.883 0.293 0.761 0.103
50 0.886 0.886 0.020 0.223 0.053 0.182 0.880 0.267 0.713 0.058
[…]
60 0.901 0.901 0.024 0.237 0.064 0.196 0.882 0.289 0.779 0.114
50 0.893 0.893 0.022 0.223 0.055 0.182 0.879 0.262 0.734 0.011
[…]
60 0.760 0.759 0.012 0.139 0.030 0.117 0.865 0.158 0.672 0.028
50 0.626 0.616 0.007 0.084 0.012 0.072 0.850 0.091 0.625 0.014
[…]
60 0.755 0.740 0.022 0.208 0.053 0.170 0.876 0.249 0.714 0.021
50 0.696 0.666 0.020 0.192 0.047 0.156 0.873 0.225 0.670 0.014
Table 2: Impact of retaining random and bottom $k\%$ tokens using the two variants of FrugalPrompt with the best performing models for each task: text classification (CLS), text summarization (SUM), question answering (QA), and text reasoning (RSN).
5 Result and Analysis
5.1 Performance across Tasks
Table 1 reveals that performance degradation becomes more pronounced as the retention threshold decreases, particularly below $60\%$ token retention, with distinct patterns emerging across tasks. For instance, at $50\%$ retention, sentiment classification accuracy drops by $5$-$10\%$ and commonsense QA accuracy by $10$-$15\%$, while summarization metrics like ROUGE-1 decline by $20$-$25\%$, as seen in Figure 2,3. The performance decline is noticeably smaller for higher values of $k$, e.g.,, GlobEnc using $80\%$ tokens achieves close to the baseline performance in classification and QA tasks.
Mathematical reasoning exhibits the steepest fall, with pass@1 scores plummeting to below $15\%$ at lower thresholds, reflecting the loss of critical numerical and logical connectors. We hypothesize that this degradation is more pronounced for reasoning tasks due to their reliance on the sequential integrity of tokens and numerical data. Low $k\%$ thresholds exacerbate this by discarding essential context, amplifying errors in tasks requiring unbroken symbolic chains, while simpler classification and summarization tasks generally tolerate greater omission by inferring from salient cues.
Model Name Llama-3 8B Llama-3 70B GPT-3.5 Gemini-2.0 FT o3-mini
Cost/1M Input Tokens (\$) 0.03 0.30 0.50 0.10 1.10
Cost/1M Output Tokens (\$) 0.06 0.40 1.50 0.40 4.40
Table 3: Cost comparison of models per 1 million tokens. The Gemini 2.0 Flash Thinking model has output cost fluctuations due to the number of thinking tokens produced.
5.2 Choice of Attribution Method
Following both Table 1 and Figure 2, for classification and QA tasks, GlobEnc outperforms DecompX across all metrics. However, in reasoning, DecompX outshines GlobEnc. DecompX’s subword-level attribution that includes linear activations may help retain more important tokens for reasoning, but might overwhelm the models in other tasks. For summarization (Figure 3), GlobEnc usually outperforms DecompX, except for the GPT-3.5 model. While DecompX is reported to produce better attribution scores Modarressi et al. (2023), most models find the token sequence filtered by word-level attention, like GlobEnc, to be more useful in retaining performance across most tasks.
5.3 Performance vs. Cost
The cost per million tokens for each model was taken from the respective provider’s website, at the time of writing, and is reported in Table 3. Note that the price is subject to change. The cost of the open-source Llama models was estimated using the lowest available rate on OpenRouter.2
Figures 2 and 3 show the performance difference between the reduced set of tokens and $100\%$ tokens for all the models. For sentiment analysis and question answering, we observe that using a reduced set of tokens generally maintains performance and shows a gentler decline as we decrease the cost. However, reducing the tokens sharply declines the performance in reasoning tasks, making it unsuitable for token reduction. The results are mixed for summarization tasks: while most models have a softer performance decline across the metrics, methods such as GPT-3.5 using GlobEnc experience rapid performance degradation.
5.4 Performance across Models
Figures 2 and 3 illustrate that for smaller models, token attribution offers negligible cost savings relative to the sharper performance degradation it entails. Consequently, token reduction is more practical for larger, costlier models such as o3-mini. Legacy models such as GPT-3.5 often display abrupt drops in performance, e.g., in summarization using GlobEnc, necessitating greater caution when reducing tokens.
5.5 Random & Bottom $k$ tokens
We select random $k\%$ and the bottom $k\%$ tokens (based on attribution scores), while preserving the text order, to establish a baseline against our token reduction method. As shown in Table 2, classification, summarization, and question answering exhibit strong retention in performance across varying thresholds. This strongly suggests potential task contamination Li and Flanigan (2024), as the ground truth answers of the test sets are easily available on the internet. Reasoning tasks experience a sharper decline, indicating genuine reliance on contextual information for solving the task.
6 Conclusion
The surging cost of large language models has often made it inaccessible to people in under-resourced environments, while increasing carbon footprint and inference time latency. Our work aims to bridge this gap by introducing prompt compression at a token-attribution level, which can reduce the cost of LLM inference and make it accessible to a wider range of users. We also reveal potential instances of task contamination that should be considered while evaluating models. We wish to explore adaptive frugalization strategies that dynamically adjust token retention based on task complexity and model confidence.
Limitations
As our token reduction pipeline heavily relies on attribution methods, misalignment between salience and true causal importance could produce suboptimal performance. The task specificity of the attribution methods also makes them unsuitable for certain tasks, such as reasoning and evaluation in multi-task settings.
We also present a few ethical considerations. As identified in our work, performance retention in certain tasks might stem from the model’s exposure to benchmark data during pretraining. This raises concerns about the validity of LLM evaluation and the need for contamination-free datasets. Prompt compression can also potentially remove safety-critical tokens and hence should be used with caution.
Appendix A Preliminaries: Token Attribution
To identify which tokens in an input text sequence carry significant influence on the language model’s decision-making process for a given task, we compute the token attribution scores using task-specific BERT-class encoders Devlin et al. (2019). We obtain the saliency scores from these models by adopting two state-of-the-art (SoTA) token attribution methods: the global token attribution analysis method (GlobEnc) Modarressi et al. (2022) and the globally propagated locally decomposed attribution analysis method (DecompX) Modarressi et al. (2023), which are discussed in this section.
A.1 GlobEnc
GlobEnc aggregates the layerwise analysis method $\mathcal{N}_{\text{{Enc}}}\coloneqq(\|\mathbf{\tilde{\text{$x$}}}_{i\leftarrow j}\|)\in\mathbb{R}^{n\times n}$ from input token $j$ to output token $i$, across all the encoder layers using attention rollout Abnar and Zuidema (2020). As elucidated in Kobayashi et al. (2020), $\tilde{x}_{i\leftarrow j}$ approximates the attributed vector contribution from input token $j$ to output token $i$, computed as the weighted and transformed value vector $\alpha_{ij}f(x_{j})$, where $f$ is the affine transformation to value space. The vector is adjusted for residual connections and layer normalization to capture the true mixed influence beyond attention weights. For each encoder layer, the output is computed as:
$\mathbf{\tilde{\text{$z$}}}_{i\leftarrow j}$ $=g_{z_{i}^{+}}\left(\sum_{h=1}^{H}\alpha_{ij}^{h}f^{h}(x_{i})+\mathbbm{1}[i=j]x_{i}\right)$ (9)
$\mathbf{\tilde{\text{$x$}}}_{i\leftarrow j}$ $\approx g_{\mathbf{\tilde{\text{$z$}}}_{i}^{+}}(\mathbf{\tilde{\text{$z$}}}_{i\leftarrow j})=\frac{\mathbf{\tilde{\text{$z$}}}_{i\leftarrow j}-\operatorname{mean}(\mathbf{\tilde{\text{$z$}}}_{i\leftarrow j})}{\operatorname{std-dev}(\mathbf{\tilde{\text{$z$}}}_{i}^{+})}\odot\gamma$ (10)
such that $\mathcal{N}_{\text{{ResLN}}}\coloneqq(\|\mathbf{\tilde{\text{$z$}}}_{i\leftarrow j}\|)\in\mathbb{R}^{n\times n}$ is the attribution analysis to measure the influence of an encoder layer’s input token $j$ on its corresponding output token $i$, considering the residual connections and layer normalizations in the attention block. Here, $\alpha_{ij}^{h}$ denotes the raw attention weight from the $i$th token to the $j$th token in the attention head $h\in\{1,\dots,H\}$, collectively forming the set of attention weights $\mathbf{A}^{h}=\{\alpha_{ij}^{h}\bigm|1\leq i,j\leq n\}$ for the $h$th attention head and $f^{h}(x)$ is the transformation function as defined in Kobayashi et al. (2020). For layer $l$, the attention rollout with respect to the input is computed as per the following recursive function:
[…]
$\hat{\mathbf{A}}_{l}$ $=0.5\bar{\mathbf{A}}_{l}+0.5\mathbf{I}$ (12)
where $\bar{\mathbf{A}}_{l}$ is the mean raw attention grid across all the heads relevant to the layer $l$.
A.2 DecompX
While GlobEnc aggregates layer-wise analyses with a modified attention rollout and approximates FFNs via residuals and scalar norms, it risks information loss. DecompX improves this by propagating the locally decomposed vectors throughout the layers to construct a global decomposition, including FFNs with linear activation approximations, and extending attributions to prediction-based values. Formally, the locally decomposed vector at layer $\ell$ is expressed as:
$\mathbf{\tilde{\text{$z$}}}^{\ell}_{i}$ $=\mathrm{LN}\left(\sum_{k=1}^{N}\left[\mathbf{x}^{\ell}_{i\Leftarrow k}+\mathbf{z}^{\ell}_{i\Leftarrow k}\right]\right),$ (13)
where the normalized contribution is given by:
$\mathrm{LN}\left(\mathbf{z}_{i}^{+\ell}\right)$ $=\sum_{k=1}^{N}\underbrace{g_{z^{+\ell}_{i}}\left(\mathbf{z}_{i\Leftarrow k}^{+\ell}\right)}_{\tilde{\mathbf{z}}_{i\Leftarrow k}^{\ell}}+\boldsymbol{\beta}$ (14)
$g_{z^{+\ell}_{i}}\left(\mathbf{z}_{i\Leftarrow k}^{+\ell}\right)$ $=\frac{\mathbf{z}_{i\Leftarrow k}^{+\ell}-m\left(\mathbf{z}_{i\Leftarrow k}^{+\ell}\right)}{s\left(\mathbf{z}_{i}^{+\ell}\right)}\odot\boldsymbol{\gamma}$
DecompX explicitly integrates all encoder layer components, notably the non-linear feed-forward networks, to enhance representational fidelity. The FFN operation is defined as:
$\mathbf{z}_{\text{FFN}}^{\ell}$ $=\text{FFN}(\tilde{\mathbf{z}}_{i}^{\ell})$ (15)
$=f_{\text{act}}(\underbrace{\tilde{\mathbf{z}}_{i}^{\ell}\mathbf{W}_{\text{FFN}}^{1}+\mathbf{b}_{\text{FFN}}^{1}}_{\bm{\zeta}_{i}^{\ell}})\mathbf{W}_{\text{FFN}}^{2}+\mathbf{b}_{\text{FFN}}^{2}$
$z_{\text{FFN},i}^{\ell}$ $=f_{\text{act}}^{(\zeta_{i}^{\ell})}\left(\sum_{k=1}^{N}\zeta_{i\Leftarrow k}^{\ell}\right)W_{\text{FFN}}^{2}+b_{\text{FFN}}^{2}$ (16)
$=\sum_{k=1}^{N}\underbrace{\theta(\zeta_{i}^{\ell})\odot\zeta_{i\Leftarrow k}^{\ell}+b_{\text{FFN}}^{2}}_{z_{\text{FFN},i\Leftarrow k}^{\ell}}$
A propagation mechanism is implemented for decomposed token representations across multiple layers, preserving their distinct identities and preventing representation conflation:
$x_{i}^{\ell+1}$ $=\operatorname{LN}\left(\sum_{k=1}^{N}\underbrace{\left[\tilde{z}_{i\Leftarrow k}^{\ell}+z_{\mathrm{FFN},i\Leftarrow k}^{\ell}\right]}_{z_{\mathrm{FFN}+,i\Leftarrow k}^{\ell}}\right)$ (17)
$=\sum_{k=1}^{N}\underbrace{gz_{\mathrm{FFN}+,i}^{\ell}\left(z_{\mathrm{FFN}+,i\Leftarrow k}^{\ell}\right)+\beta}_{x_{i\Leftarrow k}^{\ell+1}}$
The final classification mechanism utilizes these decomposed vectors to quantitatively ascertain the precise directional impact (positive or negative) exerted by each input token upon specific target classes. The established formulations within this section permit the derivation of $x_{i\Leftarrow k}^{\ell+1}$ from its antecedent, $x_{i\Leftarrow k}^{\ell}$. Through the iterative application of this procedure across the entirety of the layered architecture, the final value, $x_{i\Leftarrow k}^{L+1}$, is consequently ascertained. For the task-specific classification head,
$y_{c}=\sum_{k=1}^{N}y_{c\Leftarrow k}$ (18)
where the term $y_{c\Leftarrow k}$ quantifies the specific contribution of the $k{\text{th}}$ token towards the overall prediction score.
Appendix B Proof
We provide the derivation of Theorem 1.
Proof.
We order $D=\{d_{1},\dots,d_{q}\}$ and let $A_{r}:=\{d_{1},\dots,d_{r}\}$ be the set of tokens deleted after $r$ steps and $A_{0}=\emptyset$. For each $r$ and $u\in A_{r-1}$,
$\Delta_{d_{r},u}f(A_{r-1}\setminus\{u\})\leq\gamma_{\tau}$ [Eq. 5]
$\implies$ $\Delta_{d_{r}}\Delta_{u}f(A_{r-1}\setminus\{u\})\;\leq\;\gamma_{\tau}$ [Eq. 3]
$\implies$ $\Delta_{d_{r}}f(A_{r-1})-\Delta_{d_{r}}f(A_{r-1}\setminus\{u\})\leq+\gamma_{\tau}$
$\implies$ $\Delta_{d_{r}}f(A_{r-1})\leq\Delta_{d_{r}}f(A_{r-1}\setminus\{u\})+\gamma_{\tau}.$
Iterating over all $u\in A_{r-1}$ gives
$\Delta_{d_{r}}f(A_{r-1})$ $\leq\Delta_{d_{r}}f(\emptyset)+\gamma_{\tau}(r-1)$
$\leq f(\{d_{r}\})-f(\emptyset)+\gamma_{\tau}(r-1)$
$\leq f(\{d_{r}\})+\gamma_{\tau}(r-1)$
[…]
Formulating the telescopic sum:
$f(D)=f(A_{q})$ $=\sum_{r=1}^{q}(f(A_{r})-f(A_{r-1}))$
$=\sum_{r=1}^{q}\Delta_{d_{r}}f(A_{r-1})$ [Eq. 2]
$\leq\sum_{r=1}^{q}C_{\tau}s_{d_{r}}+\gamma_{\tau}\sum_{r=1}^{q}(r-1).$
$\text{Hence, }f(D)$ $\leq C_{\tau}\sum_{i\in D}s_{i}+\frac{\gamma_{\tau}}{2}q(q-1).$$\qed$